# Head-wise Shareable Attention for Large Language Models

### v1

@@ D.3 Visualization Study on the Shared Weights
Appendix E Generations from Models after Weight Sharing
Table E, 14 and 15 show more examples of the models after head-wise weight sharing. We present the generation results of both the shared model via DirectShare and PostShare.
@@ Appendix E Generations from Models after Weight Sharing
Llama 2-7B w/ DirectShare
The future is going to be one that presents many challenges, but also a great deal of opportunity. Everybody knows this, and it’s why we have been working so hard on the newest version of the future. The problem with the future is that it’s not always easy to see what you are doing wrong. It can be difficult to find
Llama 2-7B w/ PostShare
The future is going to be one that presents many challenges, and the world will need a new generation of leaders. The United Nations has identified the following as key issues for the future: Climate change, Food security, Energy security, Water scarcity, Global health, Human rights, International law, Peacekeeping, Sustainable
Model Generated Responses
Baichuan 2-7B w/ DirectShare (English)
There are many reasons to support why persever- ance is of great importance. The first and foremost, it’s the only way you can get better at what ever your doing! It also gives us an opportunity to learn from our mistakes or
@@
Baichuan 2-7B w/ DirectShare (Chinese)
如何健康的生活 1. 注意休息,保证睡眠。2.合理饮食:多吃水果、蔬菜等。3.适当运动(如散步) 4.保持乐观心态 5.多饮水 6.避免劳累7.不要熬夜 8.劳逸结合 9.注意安全 10.尊重生命
想要成功需要把握什么?1. 首先,要有一个好的心态。2.其次,要有一定的经济基础3.接着就是自己的实力了4. 最后一个条件是:自己一定要努力!
@@
GPT2-small w/ PostShare
In the letter, Mr Snowden said he believed the support of the intern- ational community could persuade the US government to abandon cr- iminal charges against him.<to-fr> Dans la lettre, M. le député a dit que l’appui de la communauté internationale pourrait convaincre le gouvernement américain de abandonner la charge pénale contre lui.
Dans sa lettre, M. Snowden a écrit qu’il pensait que le soutien de la communauté internationale pourrait persuader le gouvernement américain d’abandonner les charges pénales retenues contre lui.
The letter extends an offer to coop- erate with German authorities "wh- en the difficulties of this humanita- rian situation have been resolved." <to-fr> La lettre élargit une invitation à coopérer avec les autorités allemands lorsque les difficultés de cette situation humanitaire sont réglées.
La lettre avance une offre de coopération avec les autorités allemandes « lorsque les difficultés de cette situation humanitaire auront été résolues ».
The first test plane was unveiled in March and took flight for the first time in September after months of delays.<to-fr> Le premier étudiant a été démontré en mars et a fait l’avion pour la première fois après des mois de retard.
Le premier avion d’essai a été dévoilé en mars et s’est envolé pour la première fois en septembre après des mois de retard.
@@
Table 13: Responses from Llama 2-7B (Sharing Ratio=30%).

### v2

@@ F.6 Visualization Study on the Shared Weights
Appendix G More Comparison Results
G.1 Compared with Grouped Query Attention
Grouped Query Attention (GQA) has been widely adopted in LLMs to reduce the number of attention computations required, by grouping the query vectors into a smaller number of groups. Although GQA might be less flexible in choosing the reduction ratio of parameters, we also provide a performance comparison with our method. To equip the original Llama2-7B model with GQA, we perform the standard conversion proposed by  Ainslie et al. (2023) and uses neighbor grouping with group size=2 to merge heads. The performance comparison is shown in Table 14. Furthermore, we compute the average scores per category (Reasoning, NLU, Knowledge) along with the average scores across all benchmarks (Avg.) to provide a more intuitive comparison in Table 17. Overall, our method demonstrates superior performance over GQA.
Reasoning NLU Knowledge Avg.
DirectShare Ratio=10% 44.51 32.15 48.14 41.60
GQA group size=2 37.06 25.24 34.96 32.42
DirectShare Ratio=30% 44.74 28.09 39.35 37.39
Table 17: Average comparison results with GQA.
G.2 Compared with Modified Layer-wise Weight Sharing
We state the reason why choosing model pruning works as baselines in Section C.1. Prior layer-wise weight sharing techniques are proposed to apply to small-scale models like BERT. Considering one representative layer-wise weight sharing method Takase and Kiyono (2021b), we need to modify it to support LLMs, denoted as SEQ. For fair comparison, we perform layer-wise weight sharing in latter 60% layers instead of all layers. The overall performance comparison based on Llama2-7B model is presented in Table 15. Moreover, we calculate the overall average scores across all benchmarks (Avg.), the average scores for each category (Reasoning, NLU, Knowledge), and the average performance percentages relative to the original model across all benchmarks (Per.). The results in Table 18 indicate that our head-wise weight sharing method performs better than layer-wise weight sharing.
Ratio Avg. Per. Reasoning NLU Knowledge
Llama2-7B 0% 43.71 100% 44.96 35.26 50.90
DirectShare 10% 41.60 95.17% 44.51 32.15 48.14
SEQ 10% 40.74 93.21% 44.52 31.18 46.53
DirectShare 30% 37.39 85.54% 44.74 28.09 39.35
SEQ 30% 34.97 80.00% 41.68 26.32 36.90
Table 18: Average comparison results with SEQ.
Appendix H Generations from Models after Weight Sharing
Table H, 20 and 21 show more examples of the models after head-wise weight sharing. We present the generation results of both the shared model via DirectShare and PostShare.
Model Generated Responses
Llama2-7B w/ DirectShare The future is going to be one that presents many challenges, but also a great deal of opportunity. Everybody knows this, and it’s why we have been working so hard on the newest version of the future. The problem with the future is that it’s not always easy to see what you are doing wrong. It can be difficult to find
Llama2-7B w/ PostShare The future is going to be one that presents many challenges, and the world will need a new generation of leaders. The United Nations has identified the following as key issues for the future: Climate change, Food security, Energy security, Water scarcity, Global health, Human rights, International law, Peacekeeping, Sustainable
@@ Appendix H Generations from Models after Weight Sharing
Baichuan2-7B w/ DirectShare (English) There are many reasons to support why persever- ance is of great importance. The first and foremost, it’s the only way you can get better at what ever your doing! It also gives us an opportunity to learn from our mistakes or
@@
Baichuan2-7B w/ DirectShare (Chinese) {CJK*}UTF8gbsn如何健康的生活 1. 注意休息,保证睡眠。2.合理饮食:多吃水果、蔬菜等。3.适当运动(如散步) 4.保持乐观心态 5.多饮水 6.避免劳累7.不要熬夜 8.劳逸结合 9.注意安全 10.尊重生命
{CJK*}UTF8gbsn想要成功需要把握什么?1. 首先,要有一个好的心态。2.其次,要有一定的经济基础3.接着就是自己的实力了4. 最后一个条件是:自己一定要努力!
@@
GPT2-small w/ PostShare In the letter, Mr Snowden said he believed the support of the international community could persuade the US government to abandon criminal charges again- st him.<to-fr> Dans la lettre, M. le député a dit que l’appui de la communauté internationale pourrait convaincre le gouvernement américain de abandonner la charge pénale contre lui. Dans sa lettre, M. Snowden a écrit qu’il pensait que le soutien de la communauté internationale pourrait persuader le gouvernement américain d’abandonner les charges pénales retenues contre lui.
The letter extends an offer to co- operate with German authorities "when the difficulties of this hu- manitarian situation have been resolved."<to-fr> La lettre élargit une invitation à coopérer avec les autorités allemands lorsque les difficultés de cette situation humanitaire sont réglées. La lettre avance une offre de coopération avec les autorités allemandes « lorsque les difficultés de cette situation humanitaire auront été résolues ».
The first test plane was unveiled in March and took flight for the first time in September after mo- nths of delays.<to-fr> Le premier étudiant a été démontré en mars et a fait l’avion pour la première fois après des mois de retard. Le premier avion d’essai a été dévoilé en mars et s’est envolé pour la première fois en septembre après des mois de retard.